A 75kg astronaut in space against a 200 kg anvil. when the astronaut lets go of the anvil while pushing it what happens?
1
Expert's answer
2011-09-08T12:15:42-0400
Due to the Law of Momentum conservation:
mv = MV
75v = 200V, where v is the velocity of the astronaut, V is the velocity of anvil.

Thus the astronaut and anvil would move in the opposite directions and the ratio of their velocities would be v/V = 200/75 =8/3.